Operations Report – March 31st

Crew 191 Operations Report 31/03/2018

SOL:7

Name of person filing report: Wataru Okamoto

Non-nominal systems: Solar System, Trailer water

Notes on non-nominal systems:

  • – Solar system seems still unusual from Sol5. We are keeping a detailed monitoring. The system shows us P05 means the charger itself works well, but the number of SOC gave us wrong parameter. We will keep monitoring tomorrow.
  • – There are white Chemical substances on the water in trailer tank. I attach the photo.

Generator (hours run): 13hours

Solar SOC 89% (Before generator is run at night)

Diesel Almost FULL

Propane 72psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) ? 6gallons.

Water (trailer) 20%

Water (static) 100%

Trailer to Static Pump used ? no

Water (loft) Static to Loft Pump used yes

Water Meter: 132889.1
Toilet tank emptied: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) –

Oil Added –

ATV Fuel Used: N/A
Hours the ATVs were Used today: N/A

Notes on ATVs: N/A

Deimos rover used: –
Hours:116.9
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: –
Hours:27.9h
Beginning charge:100%
Ending charge: 100%
Currently charging: yes

Opportunity rover used: –
Hours: 28.4h
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Curiosity rover used: –
Hours:23.7h
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

HabCar used and why, where? –

General notes and comments: Solar system seems unusual from before yesterday. The findings of the monitoring are that solar charger system looks fine, but the number of monitor telling us wrong result for some while. We will keep monitoring. We attach the file of SOL7 solar battery parameter.

Summary of internet: Under control

Summary of suits and radios: Suits are all fine. We will make the list of radios.

Summary of Hab operations: No problem

Summary of GreenHab operations: No problem

Summary of ScienceDome operations: Only use checking the solar system.

Summary of RAMM operations: N/A

Summary of health and safety issues: No problem

Questions, concerns and requests to Mission Support: N/A

EVA Report – March 31st

EVA Report

EVA #7 – 31 Mar 2018 – Sol7
Crew members: Fumiei Morisawa (EVA leader), Kai Takeda

Objectives:
– Operational check
– Water moving from Sub tank to Main tank

Operations:
14:05 – Entering the airlock
14:08 – EVA start
14:53 – BACK
14:54 – Entering the Airlock
14:57 – Enter HAB

Narrative:
Today is day-off so today’s EVA was engineering check only. As CREW191 mission reached turnaround point, we moved water from sub-tank to main tank. In the operation we found something like glue on the surface of water in sub-tank. That looks like something chemical materials rather than biological pollution. Anyway, it may be just a little unpleasant for our health.

Vehicles
– none

Greenhab Report – March 31st

GreenHab Report
Kai Takeda
31 MAR 2018

Environmental control: (Choose which of the following is appropriate and explain further if needed)

Ambient with window/door open
Shade cloth on
Fan on (15:30~19:00)

Average temperatures: 29.8℃
Low 14.6℃
High 39.6℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 4gal

Time(s) of watering for crops: twice

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative:
The battery system was in trouble today noon.
That’s why I couldn’t switch on the fan until 15:30.

Support/supplies needed: None

Crew Photos – March 31st

SOL Essay (Photo Description)
31 Mar 2018

Sol 07
Miho Tsukishiro (XO)

Today we had a day off! So nothing to describe, instead of that, we introduce our non-human crew: two robot-dog “Aibo”s and the automatic hydroponic machine “foop”.
Aibo, the lovely dogs ease the tense atmosphere in our closed life.
The machine foop is made for interior and we can see the process of plant’s growth.
They do NOT attend briefing, clean up the HAB and of course no EVA, but they are definitely part of our team and big support for us.

Sol Summary – March 31

Crew 191 Sol Summary Report
31 Mar 2018

Sol 07

Summary Title:
A Day Off
Miho Tsukishiro (XO)

Mission Status:
in progress

Sol Activity Summary:
12:00 Lunch
14:00 EVA #7 (leader: Fumiei Morisawa, Kai Takeda) Habcom: Miho Tsukishiro
18:00 Gathering reports
19:00 CapCom

Look Ahead Plan:
RVA #8: Outreach EVA (Yusuke’s Project)

Anomalies in work:
N/A

Weather:
Partly cloudy. Almost no winds.

Crew Physical Status:
Good.

EVA:
See EVA Report.

Reports to be file:
– SOL Summary Report
– SOL Essay
– Engineer Report
– GreenHab Report
– Photo Report
– EVA Report
– EVA Request

Support Requested:
N/A